OpenStudy (anonymous):

find the integral 5^-x dx

zepdrix (zepdrix):

\(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits 5^{-x}\;dx\]

zepdrix (zepdrix):

Do you remember the derivative of this?\[\Large\bf\sf 5^x\]

OpenStudy (anonymous):

is with natural log???

zepdrix (zepdrix):

Yes :o

rvc (rvc):

5^-x/log5 +c is this correct? @zepdrix

zepdrix (zepdrix):

You weren't the person who asked the question...

rvc (rvc):

i know so i can't ans

OpenStudy (anonymous):

so is the derivative of 5^x just ln5???

rvc (rvc):

@zepdrix

OpenStudy (anonymous):

and if it were negative would it be -ln5?

zepdrix (zepdrix):

\[\Large\bf\sf (5^x)'\quad=\quad 5^x \ln 5\]

rvc (rvc):

yes

zepdrix (zepdrix):

With exponentials, you get the same thing back, but with a factor of log of the base.

zepdrix (zepdrix):

So when we integrate, instead of multiplying by that factor of ln5, we'll divide by it.

rvc (rvc):

general formula

OpenStudy (anonymous):

so ln5/5^x?

zepdrix (zepdrix):

Same thing with the chain rule. Instead of getting an extra -1 from the chain rule, we'll instead divide by a -1. If you want we can be more formal about it though :) Rewrite our integral like this,\[\Large\bf\sf \int\limits 5^{-x}\;dx\quad=\quad \int\limits e^{\log(5^{-x})}\;dx\]And then we can do a u-sub.

zepdrix (zepdrix):

Yah looks like you've got the right idea, let's just deal with the negative as well :U

rvc (rvc):

what about me @zepdrix u didnot ans me

OpenStudy (anonymous):

ooh so -ln5/5^x or no????

zepdrix (zepdrix):

Yah that looks good \c:/

zepdrix (zepdrix):

@rvc I just wish you would help instead of `posting a direct answer`. That's really frustrating.

rvc (rvc):

but was i correct? @zepdrix

OpenStudy (anonymous):

and now u substitution? what would i set u equal to??? 5^x??

zepdrix (zepdrix):

Hmm I think you missed a negative rvc :(

zepdrix (zepdrix):

No, the u-sub is a different method for solving the problem. It takes a little longer, but it's a good practice method if you're not totally comfortable with integrals yet. We don't need to u-sub, we already solved it.

rvc (rvc):

yes its -ve

rvc (rvc):

so my ans was not direct understood

rvc (rvc):

@zepdrix

OpenStudy (anonymous):

oh that wasnt too bad! alrighty thank you!!!

zepdrix (zepdrix):

np \c:/